Consider a scenario in which you could personally use present or future value calculations to come to a conclusion. Explain the scenario, and provide calculations that show how you applied the time value of money concepts.
Answer #1

Time value of money means a dollar today is more valuable than a dollar a year hence.we use rate to interest to express the time value of money.Time value of money measures the present value of future cash inflows by applying the opportunity cost of funds. Getting one dollar today is better than getting a dollar tomorrow, because money loses its value over time.

FOR EXAMPLE

If you are offered a choice between having 10000 today and having 10000 at a future date,you will usually prefer to have 10000 now.similiarly if the choice is between paying 10000 now or paying the same 10000 at future date,you will usually prefer to pay 10000 later.It is simple common sense.In the first case by accepting 10000 early ,you can simply put the money in the bank and earn some interest.similarly in the second case by deferring the payment you can earn interest by keeping the money in the bank.Therfore the time gap allowed helps us to make some money.This incremental gain is the time value of money.

Example : If 10% rate of interest and the inflation is 6%. If you buy 10,000 today and invest in a bank, the money will become 11000 one year from now. So, It is always better to get dollar today than tomorrow.

Future value = Principal + Interest amount

Invest in bank 10000 *10%

= 1000 + 10000 = 11000

In case of present value

Using the bank interest rate as the discount rate, we can find out the present value of 11000 by applying the formula.

Future cash flows/(1+i)^1 = I denotes interest rate

11000/(1+.10)^1 = 10000

The present value of future cash flow 11000 is 10000.